OpenStudy (anonymous):

Find the area of the circle with the given radius or diameter. Use = 3.14. d = 13

OpenStudy (anonymous):

r=d/2 so r=6.5 A=pi*r^2 A=132.665

OpenStudy (anonymous):

How difficult can this be, really? Memorize the following formulas and u will save yourself a lot of energy not to mention time that you could put to better use. Area of Circle is \[\pi r^2\] Circumference of Circle is \[2\pi r\] "Exact" means leave the answer in terms of pi.